# LongCodeBench: Evaluating Coding LLMs at 1M Context Windows

### v2

@@ Abstract
Context lengths for models have grown rapidly, from thousands to millions of tokens in just a few years. The extreme context sizes of modern long-context models have made it difficult to construct realistic long-context benchmarks – not only due to the cost of collecting million-context tasks but also in identifying realistic scenarios that require significant contexts. We identify code comprehension and repair as a natural testbed and challenge task for long-context models and introduce LongCodeBench (LCB), a benchmark to test LLM coding abilities in long-context scenarios. Our benchmark tests both the comprehension and repair capabilities of LCLMs in realistic and important settings by drawing from real-world GitHub issues and constructing QA (LongCodeQA) and bug fixing (LongSWE-Bench) tasks. We carefully stratify the complexity of our benchmark, enabling us to evaluate models across different scales – ranging from Qwen2.5 14B Instruct to Google’s flagship Gemini model. We find that long-context remains a weakness for all models, with performance drops such as from 29% to 3% for Claude 3.5 Sonnet, or from 70.2% to 40% for Qwen2.5.
@@ 1 Introduction
The LCB dataset is available publicly at https://huggingface.co/datasets/Steefano/LCB.
@@ 3.2 LongCodeQA
Data validation We design the data generation pipeline for LongCodeQA to satisfy two key criteria: reliability and fairness. To assess reliability, we perform manual verification on a stratified random subset of $9$ questions for each context length bracket. We verify that for $52$ of the $54$ selected questions—$96.3\%$—in-depth research about the repository and its internal codebase is required to answer successfully. Regarding fairness, a central design decision in our pipeline is to generate all questions using only the issue “Discussion”, without referencing the full codebase. This ensures that the benchmark is unbiased and does not benefit from any implicit access to repository content during construction, preserving its role as a fair evaluation tool for data construction.
@@ 4 Analysis
Table 3: Performance of different models on LongCodeQA and LongSWE-Bench across increasing context lengths (32K to 1M tokens). LongCodeQA is evaluated by answer accuracy, while LongSWE-Bench is evaluated by the percentage of solved issues. The table’s top section lists open-source models, while closed-source models appear below.
@@ 4.1 Selection of LCLMs
We evaluate diverse open- and closed-source models supporting context windows up to 1M tokens. Our selection prioritizes overall performance, maximum context length, and architectural diversity. Among closed-source models, we include GPT-4o (Transformer, 128K) (OpenAI, 2024), GPT-4.1 (Transformer, 1M), Claude 3.5 Sonnet (Transformer, 256K) (Anthropic, 2024), Gemini 2 Flash, 1.5 Pro, and 2.5 Pro (Transformer and MoE, 1M) (GeminiTeamGoogle, 2024). In the open-source category, we consider Qwen2.5-14B-Instruct (Transformer, 1M) (QwenTeam, 2024), Jamba 1.5 400B Large (hybrid Transformer-Mamba with State Space Model components and MoE, 256K) (JambaTeam, 2024), and Llama 3.1 405B Instruct and 4 Scout (Transformer, 128K-1M) (LlamaTeam, 2024). This selection underscores current performance and architectural innovations in long-context language models. We report each model’s detailed inference time and cost in Appendix C.
@@ 4.2 Performance on LongCodeQA
We report LongCodeQA results across the different context length brackets in Table 3 (left section). Most models demonstrate competitive performance at short to mid-range brackets (32K–128K), with accuracy peaking between 64K and 128K for nearly all models. For example, Llama 3.1 achieves the maximum accuracy at 64K with 72.4%, and GPT-4o reaches 76.3%, the highest result in the short-mid range. Claude 3.5 Sonnet and Jamba maintain the most solid performance, all above 69% across the lower context windows. Qwen2.5 steadily improves from 61.9% at 32K to a peak of 70.2% at 512K, showing stability in longer contexts—though this trend drops drastically at 1M, where accuracy reduces at 40.0%. Similarly, Jamba has good performance up to 128K but degrades significantly at 256K, with accuracy dropping from 72.8% to 54.2%. GPT-4.1 has the best results on the task, especially showing an improvement over its predecessor, GPT-4o, and consistency across context lengths. Llama 4 also improves over Llama 3.1, with an higher context window and better performance at 64K and 128K. Gemini 2 Flash, Gemini 1.5 Pro, and Gemini 2.5 Pro show remarkably stable behavior across all brackets up to 1M, ranging between 63–72%, preserving their accuracy in extremely long contexts.
@@ 4.3 Performance on LongSWE-Bench
We report LongSWE-Bench results across context brackets in Table 3 (right section), measuring the number of successfully resolved issues per model per bracket. As expected from the difficulty of the task—real-world bug fixing requires precise code patch—the results are significantly lower compared to LongCodeQA. Despite the overall difficulty, some trends emerge. Claude 3.5 Sonnet shows the strongest performance at shorter context lengths, solving 29% of issues at 32K. It maintains relatively high success at 64K and 128K but drops at 3% sharply at 256K. Both Gemini 2 Flash and GPT-4o peak at 32K with 11% and 10% relatively and show a gradual drop across longer contexts. On the other hand, Gemini 1.5 Pro demonstrates a consistently low performance, solving between 1% and 6% across the full range; Gemini 2.5 Pro approaches Claude 3.5 at the 32K window with 23% and maintaining a consistent rate of success until 256K, decreasing to 7% at 1M. Open-source models perform significantly worse. Jamba solves up to 3% issues at 32K but drops to 0% by 256K. Qwen2.5, Llama 3.1, Llama 4, and GPT-4.1 fail to solve any issue across all tested brackets. These results reflect both the increased challenge of generative code tasks and the current gap between open- and closed-source models in long-context code repair.
@@ A.1 Performance on the small-scale dataset before and after the first filter
Model Max. Context Length Accuracy (%) before filtering Accuracy (%) after filtering
@@ B.3 Prompt for LongCodeQA task
You are a coding expert. Your task is to analyze a GitHub repository and then answer
one question about it.
@@
Please analyze the repository text, reason through the question, and then choose
among A, B, C, D answers the correct one.
Provide your final answer in the following format:
Final Answer: <LETTER>
@@ Appendix D Comparison with SWE-Bench
Figure 5: Percentage of solved issues by Claude 3.5 Sonnet on a smaller version of LongSWE-Bench, grouped by bins of equal length range. The red line indicates a trend extracted through logistic regression, with a $p$-value of $0.005$—below the threshold ($0.01$) of statistical significance.
Figure 6: Percentage of solved issues by Claude 3.5 Sonnet on a smaller version of LongSWE-Bench in which non-oracle files are retrieved with BM25, grouped by bins of equal length range. The red line indicates a trend extracted through logistic regression with a $p$-value of $0.025$—above the threshold ($0.01$) of statistical significance.

### v3

@@ Abstract
Context lengths for models have grown rapidly, from thousands to millions of tokens in just a few years. The extreme context sizes of modern long-context models have made it difficult to construct realistic long-context benchmarks – not only due to the cost of collecting million-context tasks but also in identifying realistic scenarios that require significant contexts. We identify code comprehension and repair as a natural testbed and challenge task for long-context models and introduce LongCodeBench (LCB), a benchmark to test LLM coding abilities in long-context scenarios. Our benchmark tests both the comprehension and repair capabilities of LCLMs in realistic and important settings by drawing from real-world GitHub issues and constructing QA (LongCodeQA) and bug fixing (LongSWE-Bench) tasks. We carefully stratify the complexity of our benchmark, enabling us to evaluate models across different scales – ranging from Qwen2.5 14B Instruct to Google’s flagship Gemini model. We find that long-context remains a weakness for all models, with performance drops such as from 29% to 3% for Claude 3.5 Sonnet, or from 70.2% to 40% for Qwen2.5. The LCB dataset is available publicly at https://huggingface.co/datasets/Steefano/LCB and the codebase to replicate the work on this paper at https://github.com/Zteefano/long-code-bench.
@@ 3.2 LongCodeQA
[…]
@@ 4 Analysis
Claude 3.5 Sonnet + RAG 25.55 31.18 24.83 25.95 17.19 12.77 23 23 22 21 16 22
Table 3: Performance of different models on LongCodeQA and LongSWE-Bench across increasing context lengths (32K to 1M tokens). LongCodeQA is evaluated by answer accuracy, while LongSWE-Bench is evaluated by the percentage of solved issues. The table’s top section lists open-source models, while closed-source models appear below, and finally a RAG based model is reported in the last row.
@@ 4.1 Selection of LCLMs
We evaluate diverse open- and closed-source models supporting context windows up to 1M tokens. Our selection prioritizes overall performance, maximum context length, and architectural diversity. Among closed-source models, we include GPT-4o (Transformer, 128K) (OpenAI, 2024), GPT-4.1 (Transformer, 1M), Claude 3.5 Sonnet (Transformer, 256K) (Anthropic, 2024), Gemini 2 Flash, 1.5 Pro, and 2.5 Pro (Transformer and MoE, 1M) (GeminiTeamGoogle, 2024). In the open-source category, we consider Qwen2.5-14B-Instruct (Transformer, 1M) (QwenTeam, 2024), Jamba 1.5 400B Large (hybrid Transformer-Mamba with State Space Model components and MoE, 256K) (JambaTeam, 2024), and Llama 3.1 405B Instruct and 4 Scout (Transformer, 128K-1M) (LlamaTeam, 2024). This selection underscores current performance and architectural innovations in long-context language models.
In this work, we deliberately focus on long-context evaluation rather than RAG optimization. We view these as distinct approaches to handling large documents and LongCodeBench aims to test pure long-context capabilities. For reference, we report results for Claude 3.5 Sonnet with RAG. We remark that performance of RAG architectures depends on prompt structure and retrieval techniques and therefore the scores reported here are a narrow view of RAG capabilities. We report each model’s detailed inference time and cost in Appendix C.
@@ 4.2 Performance on LongCodeQA
We report LongCodeQA results across the different context length brackets in Table 3 (left section). Most models demonstrate competitive performance at short to mid-range brackets (32K–128K), with accuracy peaking between 64K and 128K for nearly all models. For example, Llama 3.1 achieves the maximum accuracy at 64K with 72.4%, and GPT-4o reaches 76.3%, the highest result in the short-mid range. Claude 3.5 Sonnet and Jamba maintain the most solid performance, all above 69% across the lower context windows. Qwen2.5 steadily improves from 61.9% at 32K to a peak of 70.2% at 512K, showing stability in longer contexts—though this trend drops drastically at 1M, where accuracy reduces at 40.0%. Similarly, Jamba has good performance up to 128K but degrades significantly at 256K, with accuracy dropping from 72.8% to 54.2%. GPT-4.1 has the best results on the task, especially showing an improvement over its predecessor, GPT-4o, and consistency across context lengths. Llama 4 also improves over Llama 3.1, with an higher context window and better performance at 64K and 128K. Gemini 2 Flash, Gemini 1.5 Pro, and Gemini 2.5 Pro show remarkably stable behavior across all brackets up to 1M, ranging between 63–72%, preserving their accuracy in extremely long contexts. We observe that RAG negatively impacts the performance of Claude 3.5 Sonnet, achieving a score comparable to random guessing. This is an interesting phenomenon that shows the sensitivity of RAG algorithms to different tasks.
@@ 4.3 Performance on LongSWE-Bench
We report LongSWE-Bench results across context brackets in Table 3 (right section), measuring the number of successfully resolved issues per model per bracket. As expected from the difficulty of the task—real-world bug fixing requires precise code patch—the results are significantly lower compared to LongCodeQA. Despite the overall difficulty, some trends emerge. Claude 3.5 Sonnet shows the strongest performance at shorter context lengths, solving 29% of issues at 32K. It maintains relatively high success at 64K and 128K but drops at 3% sharply at 256K. Both Gemini 2 Flash and GPT-4o peak at 32K with 11% and 10% relatively and show a gradual drop across longer contexts. On the other hand, Gemini 1.5 Pro demonstrates a consistently low performance, solving between 1% and 6% across the full range; Gemini 2.5 Pro approaches Claude 3.5 at the 32K window with 23% and maintaining a consistent rate of success until 256K, decreasing to 7% at 1M. Open-source models perform significantly worse. Jamba solves up to 3% issues at 32K but drops to 0% by 256K. Qwen2.5, Llama 3.1, Llama 4, and GPT-4.1 fail to solve any issue across all tested brackets. Contrary to the LongCodeQA task, RAG helps Claude 3.5 Sonnet achieve positive scores on LongSWE-Bench, stressing again the importance of RAG strategy dependent on the deployment task.These results reflect both the increased challenge of generative code tasks and the current gap between open- and closed-source models in long-context code repair.
@@ 5 Conclusion
Acknowledgments
We thank Luca Rigazio and Yoshikuni Sato for their valuable feedback and support during the development of LongCodeBench. We also acknowledge CINECA for providing computational resources and infrastructure support—through access to the Leonardo supercomputer—that made the large-scale evaluations possible. Finally, we thank the anonymous reviewers for their constructive feedback that helped improve the quality of this work.
@@ A.1 Performance on the small-scale dataset before and after the first filter
Model Max. Context Length
Accuracy (%)
before filtering
Accuracy (%)
after filtering
@@ B.3 Prompt for LongCodeQA task
You are going to be provided the content of a repository and a question about it.
Provide the answer to the question by stating ONLY the letter associated to the
question.
@@ Appendix D Comparison with SWE-Bench
Figure 5: Percentage of solved issues by Claude 3.5 Sonnet on a smaller version of LongSWE-Bench, grouped by bins of equal length range. The red line indicates a trend extracted through logistic regression, with a $p$-value of 0.0050.005—below the threshold ($0.01$) of statistical significance.
Figure 6: Percentage of solved issues by Claude 3.5 Sonnet on a smaller version of LongSWE-Bench in which non-oracle files are retrieved with BM25, grouped by bins of equal length range. The red line indicates a trend extracted through logistic regression with a $p$-value of 0.0250.025—above the threshold ($0.01$) of statistical significance.